Solve fo x.
X^4+61x^2+900=0


Answer by htmentor(1343)

Solve for x.
X^4+61x^2+900=0
=======================
Let u=x^2
u^2 + 61u + 900 = 0
This can be factored as (u+36)(u+25) = 0
The solutions are u=-36, u=-25
x^2 = -36 -> x = 6i, -6i
x^2 = -25 -> x = 5i, -5i